# Conversational LLMs Simplify Secure Clinical Data Access, Understanding, and Analysis

### v1

@@ 4.5 Discussion of Results
The evaluation results are promising and validate the core architecture of the M3 system. The 94% accuracy on answerable questions demonstrates that a state-of-the-art large language model like Claude Sonnet 4, when provided with the proper tools and context via the MCP, can effectively query a complex, specialized database like MIMIC-IV without task-specific fine-tuning.
The qualitative analysis underscores a primary remaining challenge that is not merely technical (i.e., generating valid SQL), but semantic: correctly interpreting user intent. The error caused by linguistic ambiguity suggests that future work should focus on developing strategies for ambiguity detection and resolution. This could involve prompting the model to ask clarifying questions when it detects ambiguity or incorporating a feedback loop where users can refine the system’s interpretation.
@@ 5.1 Conclusion
At the same time, we acknowledge that M3 is only a starting point. Its current focus on MIMIC-IV, dependence on LLM quality, and narrow focus on data retrieval—not inference—highlight opportunities for deeper integration with broader research and clinical workflows. Nonetheless, the successful deployment of M3 affirms that such interfaces can meaningfully reduce friction in data exploration, and we hope this work inspires continued development and community-driven extension.
@@ Acknowledgments
The authors would like to thank Dr. Gloria Hyunjung Kwak who provided clinical domain expertise, advised on cohort definitions and validation protocols. The author PM acknowledges financial support from the Fulbright Scholarship and Erasmus Mundus JM Scholarship. M3 operates exclusively on the de-identified MIMIC-IV corpus released through PhysioNet under the standard Data Use Agreement. All investigators completed the required CITI “Data or Specimens Only” training, and access to the full BigQuery backend was provisioned inside credential-bound, read-only Google Cloud projects. No new patient-level data were collected, stored, or exported. Lowering the technical barrier to interrogating critical-care records carries non-trivial misuse risks. A malicious or careless user could: (i) attempt linkage attacks that re-identify individuals, (ii) over-interpret associative findings as causal and deploy them for bedside decision-making, or (iii) use the interface to generate incomplete cohorts. PhysioNet and M3 mitigate these threats through user training and with the safeguards discussed in the methodology 3.5. All empirical results, architectural diagrams, and dataset statistics cited in this manuscript were obtained from publicly available resources: the MIMIC-IV repository on PhysioNet, the EHRSQL-2024 benchmark materials, and the open-source M3 codebase on GitHub.

### v3

@@ 4.5 Discussion of Results
The evaluation results offer encouraging preliminary validation of the M3 architecture. The 93-94% accuracy demonstrates that language models, when provided with proper tools via MCP, can effectively query complex databases like MIMIC-IV without task-specific fine-tuning.
The near-equivalent performance of gpt-oss-20B has important implications. For researchers facing data privacy constraints, regulatory requirements, or limited connectivity, local deployment offers a practical path forward. The smaller parameter count contributes to lower computational requirements and faster inference on local hardware.
The qualitative analysis underscores that the primary challenge is semantic: correctly interpreting user intent.
The qualitative analysis underscores a primary remaining challenge that is not merely technical (i.e., generating valid SQL), but semantic: correctly interpreting user intent. Errors caused by linguistic ambiguity suggest future work should focus on ambiguity detection and resolution [min2020ambigqa]. The fact that both models encountered similar obstacles suggests these represent general challenges for natural language-to-SQL in clinical contexts [singhal2023large].
Most errors could potentially be resolved through multi-turn conversations where user and system iteratively refine query specification. For basic exploratory data analysis, M3 performs remarkably well. For complex analytical tasks involving nuanced temporal relationships, expert involvement remains valuable.
The transparency of M3’s approach, exposing generated SQL alongside natural language results, enables expert oversight. Researchers can verify that interpretations match intent and spot logical errors, maintaining analytical rigor.
@@
5 Ethical Considerations
The development and deployment of AI systems like M3 occur within societies characterized by profound forms of social, material, and political inequality. The healthcare and medical research domains are particularly susceptible to these inequalities, making it essential to address the ethical implications of technologies that democratize access to clinical data analysis [birhane2021algorithmic].
5.1 Benefits and Maintaining Analytical Rigor
M3 offers significant potential benefits by making clinical data more accessible to researchers, including those with limited computational resources or laboratory infrastructure for complex data analyses. This accessibility could advance medical knowledge and help address system-level inequalities by enabling broader participation in clinical research. The democratization of clinical data analysis represents an opportunity to engage diverse perspectives in medical research while maintaining appropriate safeguards.
The extensive training and experience previously required to conduct database queries traditionally enabled research scientists and clinicians to evaluate the scientific validity of their analyses, identify potential misinterpretations of statistical results, and understand the complexities of translating query results into clinical practice [singhal2023large]. Experienced researchers typically possess intimate knowledge of the datasets they work with, including understanding how specific categories were constructed, how data was collected, and the implications of these factors for specific research queries.
To ensure M3 users can maintain this level of analytical rigor, we recommend implementing comprehensive training programs that bridge the gap between technical accessibility and domain expertise. This includes providing detailed documentation about dataset construction, establishing mentorship programs pairing experienced researchers with new users, and creating educational resources that emphasize the importance of contextual understanding in clinical data analysis.
M3’s design principle of exposing underlying SQL queries alongside natural language results directly supports this goal by enabling users to understand and validate the analytical approach, fostering transparency and reproducibility in research workflows.
5.2 Promoting Equity and Addressing Bias
The past decade has witnessed extensive focus on bias in artificial intelligence systems, with AI algorithms shown to replicate and amplify existing forms of societal inequality and discrimination [birhane2021algorithmic]. These concerns are particularly acute in healthcare, where biased algorithms can perpetuate historical injustices and disproportionately affect communities already experiencing significant social and health inequalities [obermeyer2019dissecting].
M3 users should be equipped with training and tools to identify how specific populations, particularly marginalized groups, may be represented in datasets, and how to conduct analyses that account for potential biases. To support this, we recommend developing bias awareness training, implementing tools that help users understand the demographic composition of their analyses, and establishing review processes that evaluate research for potential equity implications.
The transparency provided by M3’s query exposition enables peer review and validation of analytical approaches, supporting the identification and correction of potential biases in research design and interpretation.
5.3 Security, Privacy, and Accountability
M3 implements comprehensive safeguards to address privacy and security concerns inherent in clinical data analysis. While MIMIC-IV consists of de-identified data, M3 maintains robust protections through comprehensive security measures. M3’s security framework includes OAuth2 authentication, query validation to prevent unauthorized operations, comprehensive audit logging, and rate limiting to prevent system abuse.
To ensure accountability in AI-assisted research, we propose a collaborative responsibility model where system developers maintain robust security measures and clear documentation, users employ the system appropriately with proper training, institutions establish governance frameworks, and the research community maintains quality standards through peer review.
The linguistic ambiguity challenges identified in our evaluation results (Section 4) highlight the importance of verification procedures. M3’s query transparency features enable users to validate their analyses and support reproducible research practices. We recommend that institutions establish procedures for reviewing AI-generated analyses, particularly those intended for clinical application or publication.
5.4 Implementation and Best Practices
To ensure M3 enhances rather than compromises scientific rigor, we recommend several best practices. Users should validate results through multiple approaches where possible, thoroughly document their analytical procedures including the natural language queries used, and ensure appropriate peer review of their work. M3’s transparency features, including exposed SQL queries and comprehensive logging, directly support these practices.
Training programs should emphasize the importance of critical evaluation skills and help users understand both the capabilities and limitations of AI-assisted analysis. By combining M3’s accessibility with robust educational frameworks, we can democratize clinical data analysis while maintaining the highest standards of scientific excellence.
Based on these considerations, we recommend institutions adopting M3 implement phased deployment strategies, beginning with supervised use in educational settings. Comprehensive training programs should address both technical and ethical aspects of AI-assisted clinical data analysis. Clear governance frameworks should establish policies for M3 usage, including guidelines for result interpretation and approval processes for sensitive analyses.
Regular monitoring of M3 usage patterns and outcomes can help identify areas for improvement and ensure alignment with institutional and professional standards. Engaging diverse stakeholders, including M3 users, clinical experts, and ethicists, will help ensure ongoing alignment with evolving best practices.
M3 represents a significant opportunity to democratize clinical data analysis while maintaining the rigor essential for advancing medical knowledge. Through careful attention to ethical considerations, comprehensive training, and robust governance frameworks, we can harness the benefits of this technology while preserving the integrity of medical research and promoting equitable access to clinical insights.
@@ 6.1 Conclusion
At the same time, we acknowledge that M3 is only a starting point. Its current focus on MIMIC-IV, dependence on LLM quality, and narrow focus on data retrieval highlight opportunities for deeper integration with broader research and clinical workflows. Nonetheless, the successful deployment of M3 affirms that such interfaces can meaningfully reduce friction in data exploration, and we hope this work inspires continued development and community-driven extension.
@@ Acknowledgments
The authors would like to thank Dr. Gloria Hyunjung Kwak who provided clinical domain expertise, advised on cohort definitions and validation protocols. The author PM acknowledges financial support from the Fulbright Scholarship and Erasmus Mundus JM Scholarship. LAC is funded by the National Institute of Health through DS-I Africa U54 TW012043-01 and Bridge2AI OT2OD032701, the National Science Foundation through ITEST #2148451, and a grant of the Korea Health Technology R&D Project through the Korea Health Industry Development Institute (KHIDI), funded by the Ministry of Health & Welfare, Republic of Korea (grant number: RS-2024-00403047). This research was supported by a grant of the Korea Health Technology R&D Project through the Korea Health Industry Development Institute (KHIDI), funded by the Ministry of Health & Welfare, Republic of Korea (grant number: RS-2024-00439677).
M3 operates exclusively on the de‑identified MIMIC‑IV corpus released through PhysioNet under the standard Data Use Agreement. All investigators completed the required CITI “Data or Specimens Only” training, and access to the full BigQuery backend was provisioned inside credential‑bound, read‑only Google Cloud projects. No new patient‑level data were collected, stored, or exported. Lowering the technical barrier to interrogating critical‑care records carries non‑trivial misuse risks. A malicious or careless user could: (i) attempt linkage attacks that re‑identify individuals, (ii) over‑interpret associative findings as causal and deploy them for bedside decision‑making, or (iii) use the interface to generate incomplete cohorts. PhysioNet and M3 mitigate these threats through user training and with the safeguards discussed in the methodology 3.5. All empirical results, architectural diagrams, and dataset statistics cited in this manuscript were obtained from publicly available resources: the MIMIC‑IV repository on PhysioNet, the EHRSQL‑2024 benchmark materials, and the open‑source M3 codebase on GitHub.
@@
CRediT Author Statement
Rafi Al Attrach: Investigation (lead), Software (lead), Writing – original draft (lead), Writing – review & editing (equal).
Pedro Moreira: Investigation (lead), Software (lead), Writing – original draft (lead), Funding acquisition (supporting), Writing – review & editing (equal).
Rajna Fani: Investigation (lead), Software (lead), Writing – original draft (lead), Writing – review & editing (equal).
Renato Umeton: Conceptualization (equal), Methodology (supporting), Investigation (supporting), Writing – review & editing (equal).
Amelia Fiske: Conceptualization (supporting), Writing – review & editing (equal), Ethics (lead).
Leo A. Celi: Conceptualization (equal), Supervision (lead), Project administration (lead), Writing – review & editing (equal).
This statement is based on CRediT, the ANSI/NISO Contributor Role Taxonomy.
Supporting Information
Detailed benchmark results including all natural language queries, model answers, gold standard answers, SQL queries, and error analysis notes are publicly available in the M3 GitHub repository: